OpenStudy (anonymous):

Solve for r. T= P+3r /6

OpenStudy (anonymous):

T-P=3r/6 6(T-P)=3r \[r=2(T-P)\]

OpenStudy (anonymous):

Can you tell me step by step in words?

OpenStudy (anonymous):

try making r the subject first take P to the other side of the equation T-P=3r/6 to get rid of 6 multiplr by 6 on both sides 6T-6P=3r to only remain with r divide by 3 all through 6T-6P=3r --------- 3 2T-2P=r since 2 is common you factor it out = 2(T-P) answer is 2(T-P)=r